# GME: Improving Universal Multimodal Retrieval by Multimodal LLMs

### v1

@@ Appendix B Results Details
| Type | Task | Dataset | VISTA | CLIP-SF | One-Peace | DSE | E5-V | GME-2B | GME-7B | ||
TT | ArguAna | 63.61 | 52.45 | 32.93 | 53.46 | 54.28 | 61.52 | 64.60 | ||
| Climate-FEVER | 31.17 | 20.00 | 20.27 | 19.79 | 21.64 | 42.30 | 45.38 | ||||
| CQADupStack | 42.35 | 30.61 | 41.32 | 36.51 | 41.69 | 38.13 | 41.66 | ||||
| DBPedia | 40.77 | 26.37 | 32.43 | 40.75 | 38.78 | 46.31 | 50.78 | ||||
| FEVER | 86.29 | 50.58 | 51.91 | 80.12 | 78.99 | 92.60 | 94.27 | ||||
| FiQA2018 | 40.65 | 22.14 | 36.79 | 36.2 | 45.41 | 45.30 | 57.14 | ||||
| HotpotQA | 72.6 | 41.33 | 46.51 | 70.79 | 60.88 | 72.93 | 79.21 | ||||
| MSMARCO | 41.35 | 22.15 | 36.55 | 37.73 | 41.23 | 40.88 | 42.38 | ||||
| NFCorpus | 37.39 | 27.05 | 31.6 | 32.82 | 36.97 | 37.20 | 38.40 | ||||
| NQ | 54.15 | 25.45 | 42.87 | 52.97 | 51.58 | 60.01 | 67.04 | ||||
| Quora | 88.90 | 81.63 | 87.46 | 85.84 | 87.6 | 87.24 | 88.05 | ||||
| SCIDOCS | 21.73 | 14.75 | 21.64 | 15.66 | 22.36 | 23.17 | 27.38 | ||||
| SciFact | 74.04 | 55.98 | 64.51 | 68.97 | 72.75 | 63.82 | 62.31 | ||||
| Touche2020 | 25.7 | 17.47 | 16.90 | 14.50 | 21.61 | 29.06 | 23.26 | ||||
| TRECCOVID | 77.90 | 63.61 | 69.28 | 52.98 | 72.85 | 59.24 | 52.60 | ||||
| WebQA | 83.80 | 84.44 | 63.67 | 83.95 | 89.94 | 95.19 | 95.85 | ||||
| II | Nights | 24.43 | 31.42 | 31.27 | 27.36 | 27.92 | 29.86 | 31.89 | |||
TI | VisualNews | 5.77 | 42.80 | 48.95 | 14.12 | 29.46 | 38.85 | 46.54 | ||
| Fashion200k | 3.08 | 18.38 | 32.34 | 3.08 | 3.78 | 25.83 | 31.82 | ||||
| MSCOCO | 47.97 | 80.75 | 71.45 | 74.62 | 52.38 | 71.82 | 75.14 | ||||
| Flickr30k | 74.68 | 94.28 | 92.78 | 94.42 | 77.38 | 92.92 | 91.92 | ||||
| TVD | TAT-DQA | 2.05 | 5.49 | 14.44 | 49.01 | 9.08 | 71.05 | 76.12 | |||
| ArxivQA | 10.30 | 24.10 | 43.94 | 78.17 | 41.16 | 83.91 | 87.58 | ||||
| DocVQA | 8.01 | 11.80 | 23.48 | 45.83 | 24.37 | 54.57 | 56.63 | ||||
| InfoVQA | 30.02 | 48.78 | 59.97 | 82.06 | 49.5 | 91.11 | 92.39 | ||||
| Shift Project | 3.26 | 6.06 | 17.02 | 69.84 | 13.16 | 94.29 | 97.26 | ||||
| Artificial Intelligence | 7.34 | 28.64 | 45.41 | 96.88 | 46.18 | 99.02 | 99.63 | ||||
| Government Reports | 6.90 | 34.67 | 55.98 | 92.04 | 53.05 | 98.89 | 99.63 | ||||
| Energy | 11.05 | 27.19 | 53.21 | 92.62 | 56.77 | 93.15 | 95.89 | ||||
| TabFQuad | 13.08 | 21.53 | 57.05 | 79.29 | 58.22 | 94.61 | 94.58 | ||||
| Healthcare Industry | 9.39 | 32.64 | 59.55 | 96.35 | 59.61 | 98.89 | 99.63 | ||||
| IT | VisualNews | 2.79 | 42.67 | 47.27 | 8.74 | 29.54 | 38.32 | 45.54 | |||
| Fashion200k | 4.72 | 18.10 | 30.89 | 3.91 | 4.62 | 27.57 | 34.20 | ||||
| MSCOCO | 48.92 | 91.94 | 85.6 | 82.06 | 86.4 | 84.12 | 84.90 | ||||
| Flickr30k | 68.50 | 99.11 | 98.60 | 97.11 | 89.62 | 97.70 | 98.70 | ||||
TIT | WebQA | 54.84 | 78.42 | 32.42 | 66.99 | 49.62 | 83.15 | 84.59 | ||
| EDIS | 36.78 | 54.09 | 53.01 | 41.26 | 49.62 | 70.32 | 77.29 | ||||
| ITT | OVEN | 22.32 | 45.98 | 23.69 | 0.38 | 14.4 | 58.17 | 63.41 | |||
| INFOSEEK | 18.53 | 27.58 | 20.05 | 3.06 | 12.69 | 39.06 | 43.14 | ||||
| ReMuQ | 76.20 | 83.71 | 26.41 | 94.60 | 52.15 | 98.75 | 98.12 | ||||
| OKVQA | 17.14 | 17.44 | 9.67 | 13.28 | 16.71 | 28.38 | 27.03 | ||||
| LLaVA | 72.81 | 91.91 | 51.64 | 53.18 | 77.48 | 98.54 | 99.18 | ||||
| ITI | FashionIQ | 3.28 | 24.54 | 2.93 | 9.81 | 3.73 | 27.50 | 31.43 | |||
| CIRR | 14.65 | 45.25 | 10.53 | 36.52 | 13.19 | 46.83 | 53.69 | ||||
| ITIT | OVEN | 27.77 | 68.83 | 30.56 | 0.39 | 54.46 | 75.98 | 80.30 | |||
| EVQA | 28.75 | 40.08 | 16.64 | 15.34 | 26.39 | 79.21 | 81.75 | ||||
| INFOSEEK | 22.27 | 49.05 | 23.32 | 5.96 | 39.69 | 44.21 | 58.80 | ||||
| Avg. | 37.32 | 43.66 | 42.01 | 50.04 | 42.52 | 64.45 | 67.44 |
B.1 Detailed Results on UMRB
@@
Table 7 presents the detailed evaluation results of the baseline systems alongside our GME on UMRB tasks. First, focusing on the average scores, our smaller model, i.e., GME-Qwen2-VL-2B, already outperforms the previous state-of-the-art UMR model (VISTA). The larger model, i.e., GME-Qwen2-VL-7B, further enhances this performance. In addition, focusing on specific scores on different datasets, our GME achieves state-of-the-art performance on each dataset except the Nights dataset. VISTA and CLIP-SF scored highly on the Nights dataset, likely due to their use of independent image and text encoders for cross-modal retrieval. In the II task, these models relied solely on the image encoder for encoding without cross-modal alignment, which may explain their superior performance on the Nights dataset.
| Type | Task | Dataset | TT | II | TVD | TI | ITIT | Mix | ||
TT | Arguan | 56.25 | 43.51 | 56.73 | 33.53 | 53.22 | 56.22 | ||
| NFCorpus | 35.23 | 28.89 | 33.23 | 33.18 | 30.48 | 35.76 | ||||
| Quora | 87.82 | 74.37 | 86.32 | 86.43 | 85.2 | 87.4 | ||||
| SCIDOCS | 19.07 | 11.82 | 17.51 | 17.2 | 16.93 | 19.88 | ||||
| TRECCOVID | 75.57 | 47.89 | 50.89 | 72.37 | 58.92 | 76.38 | ||||
| II | Nights | 27.97 | 28.11 | 24.9 | 28.53 | 26.04 | 30.85 | |||
TI | MSCOCO | 59.7 | 59.41 | 63.67 | 76.91 | 44.97 | 75.3 | ||
| Flickr30k | 83.92 | 65.52 | 87.32 | 93.18 | 74.52 | 93.06 | ||||
| TVD | DocVQA | 35.8 | 24.24 | 48.38 | 40.58 | 28.05 | 45.62 | |||
| Shift Project | 57.86 | 45.47 | 77.08 | 50.36 | 53.12 | 74.84 | ||||
| IT | MSCOCO | 74.72 | 63.82 | 80.46 | 84.64 | 70.48 | 84.24 | |||
| Flickr30k | 94.1 | 82.5 | 96.3 | 97.2 | 90.1 | 97.5 | ||||
ITT | LLaVA | 92.75 | 89.05 | 86.02 | 89.24 | 88.73 | 95.02 | ||
| ReMuQ | 89.61 | 85.47 | 76.45 | 85.12 | 86.73 | 89.75 | ||||
| OKVQA | 24.55 | 16.6 | 15.78 | 16.92 | 18.57 | 20.23 | ||||
| ITI | FashionIQ | 5.53 | 4.2 | 5.43 | 8.86 | 11.08 | 11.89 | |||
| CIRR | 17.24 | 15.04 | 15.42 | 17.5 | 25.71 | 29.86 | ||||
| ITIT | OVEN | 59.81 | 38.42 | 57.31 | 56.69 | 65.08 | 63.04 | |||
@@
B.2 Detailed Results on UMRB-Partial
Figure 3 illustrates our exploration of the training data, as discussed in Section 4.2, with specific results presented in Table 8. This table details the scores of our models trained on six data types: TT, II, TVD, TI, ITIT, and Mix across various tasks. We find that the model trained on mixed data performs the best.
@@ B.2 Detailed Results on UMRB-Partial
| GME-Qwen2VL-2B | 53.31 | 61.52 | 42.30 | 38.13 | 46.31 | 92.60 | 45.30 | 72.93 | 40.88 | 37.20 | 60.01 | 87.24 | 23.17 | 63.82 | 29.06 | 59.24 | ||||||||||||||||||||||||||
| GME-Qwen2VL-7B | 55.63 | 64.60 | 45.38 | 41.66 | 50.78 | 94.27 | 57.14 | 79.21 | 42.38 | 38.40 | 67.04 | 88.05 | 27.38 | 62.31 | 23.26 | 52.60 | ||||||||||||||||||||||||||
B.3 Detailed Results on BEIR
@@
| GME-Qwen2-VL-2B | 53.55 | 38.85 | 71.82 | 25.83 | 95.19 | 70.32 | 83.15 | 38.32 | 84.12 | 27.57 | 29.86 | 58.17 | 39.06 | 27.50 | 46.83 | 75.98 | 44.21 | |||||||||||||||||||||
| GME-Qwen2-VL-7B | 58.66 | 46.54 | 75.14 | 31.82 | 95.85 | 77.29 | 84.59 | 45.54 | 84.90 | 34.20 | 31.89 | 63.41 | 43.14 | 31.43 | 53.69 | 80.30 | 58.80 | |||||||||||||||||||||
B.4 Detailed Results on M-BEIR
M-BEIR, the Multimodal Benchmark for IR, serves as a comprehensive large-scale retrieval benchmark designed to evaluate multimodal retrieval models. As shown in Table 10, we report Recall@10 scores for the Fashion200K and FashionIQ datasets, while Recall@5 scores are provided for all other datasets. In M-BEIR, our GME continues to demonstrate state-of-the-art performance, underscoring the effectiveness of our approach.
@@ B.4 Detailed Results on M-BEIR
| GME-Qwen2-VL-2B | 83.91 | 54.57 | 91.11 | 94.61 | 71.05 | 94.29 | 99.02 | 93.15 | 98.89 | 98.89 | 87.84 |
| GME-Qwen2-VL-7B | 87.58 | 56.63 | 92.39 | 94.58 | 76.12 | 97.26 | 99.63 | 95.89 | 99.63 | 99.63 | 89.92 |
B.5 Detailed Results on ViDoRe
ViDoRe represents the Visual Document Retrieval Benchmark, encompassing various page-level screenshot retrieval tasks. This benchmark includes the TVD tasks within our UMRB. Table 11 presents the detailed nDCG@5 scores for our GME and other models. Our smaller model, i.e., GME-Qwen2-VL-2B, surpasses the previous state-of-the-art model (ColPali), which was exclusively trained on this dataset for this specific task. The larger model, i.e., GME-Qwen2-VL-7B, further improves upon this performance.
Appendix C Experiment Details
@@
C.1 Training Details
Our GME models (both 2B and 7B) are initialized using the Qwen2-VL (Wang et al., 2024b) model series. We employ the transformers library for training in BF16 precision. The training utilizes Low-Rank Adaptation (LoRA) (Hu et al., 2022) with a rank of 8. We apply a decoupled AdamW optimizer with a learning rate and a weight decay of 1e-4. Additional hyperparameters are detailed in Table 12.

### v2

@@ 8 Results Details
| Type | Task | Dataset | VISTA | CLIP-SF | One-Peace | DSE | E5-V | GME-2B | GME-7B |
| Single- Modal | TT | ArguAna | 63.61 | 52.45 | 32.93 | 53.46 | 54.28 | 63.18 | 72.11 |
| Climate-FEVER | 31.17 | 20.00 | 20.27 | 19.79 | 21.64 | 41.08 | 48.36 | ||
| CQADupStack | 42.35 | 30.61 | 41.32 | 36.51 | 41.69 | 39.06 | 42.16 | ||
| DBPedia | 40.77 | 26.37 | 32.43 | 40.75 | 38.78 | 41.00 | 46.30 | ||
| FEVER | 86.29 | 50.58 | 51.91 | 80.12 | 78.99 | 92.06 | 93.81 | ||
| FiQA2018 | 40.65 | 22.14 | 36.79 | 36.2 | 45.41 | 43.8 | 63.23 | ||
| HotpotQA | 72.6 | 41.33 | 46.51 | 70.79 | 60.88 | 65.3 | 68.18 | ||
| MSMARCO | 41.35 | 22.15 | 36.55 | 37.73 | 41.23 | 40.61 | 42.93 | ||
| NFCorpus | 37.39 | 27.05 | 31.6 | 32.82 | 36.97 | 38.84 | 36.95 | ||
| NQ | 54.15 | 25.45 | 42.87 | 52.97 | 51.58 | 54.52 | 56.08 | ||
| Quora | 88.90 | 81.63 | 87.46 | 85.84 | 87.6 | 88.12 | 89.67 | ||
| SCIDOCS | 21.73 | 14.75 | 21.64 | 15.66 | 22.36 | 22.94 | 26.35 | ||
| SciFact | 74.04 | 55.98 | 64.51 | 68.97 | 72.75 | 74.19 | 82.43 | ||
| Touche2020 | 25.7 | 17.47 | 16.90 | 14.50 | 21.61 | 26.57 | 22.55 | ||
| TRECCOVID | 77.90 | 63.61 | 69.28 | 52.98 | 72.85 | 71.73 | 77.49 | ||
| WebQA | 83.80 | 84.44 | 63.67 | 83.95 | 89.94 | 94.34 | 94.34 | ||
| II | Nights | 24.43 | 31.42 | 31.27 | 27.36 | 27.92 | 30.61 | 30.57 | |
| Cross- Modal | TI | VisualNews | 5.77 | 42.80 | 48.95 | 14.12 | 29.46 | 39.20 | 46.27 |
| Fashion200k | 3.08 | 18.38 | 32.34 | 3.08 | 3.78 | 23.50 | 27.64 | ||
| MSCOCO | 47.97 | 80.75 | 71.45 | 74.62 | 52.38 | 76.22 | 79.77 | ||
| Flickr30k | 74.68 | 94.28 | 92.78 | 94.42 | 77.38 | 94.5 | 97.38 | ||
| TVD | TAT-DQA | 2.05 | 5.49 | 14.44 | 49.01 | 9.08 | 57.88 | 64.06 | |
| ArxivQA | 10.30 | 24.10 | 43.94 | 78.17 | 41.16 | 81.41 | 82.55 | ||
| DocVQA | 8.01 | 11.80 | 23.48 | 45.83 | 24.37 | 46.86 | 49.34 | ||
| InfoVQA | 30.02 | 48.78 | 59.97 | 82.06 | 49.5 | 84.97 | 88.79 | ||
| Shift Project | 3.26 | 6.06 | 17.02 | 69.84 | 13.16 | 77.94 | 83.5 | ||
| Artificial Intelligence | 7.34 | 28.64 | 45.41 | 96.88 | 46.18 | 95.75 | 98.02 | ||
| Government Reports | 6.90 | 34.67 | 55.98 | 92.04 | 53.05 | 92.05 | 94.05 | ||
| Healthcare Industry | 9.39 | 32.64 | 59.55 | 96.35 | 59.61 | 96.08 | 97.29 | ||
| Energy | 11.05 | 27.19 | 53.21 | 92.62 | 56.77 | 89.17 | 93.09 | ||
| TabFQuad | 13.08 | 21.53 | 57.05 | 79.29 | 58.22 | 91.79 | 94.92 | ||
| IT | VisualNews | 2.79 | 42.67 | 47.27 | 8.74 | 29.54 | 38.21 | 47.16 | |
| Fashion200k | 4.72 | 18.10 | 30.89 | 3.91 | 4.62 | 26.61 | 31.05 | ||
| MSCOCO | 48.92 | 91.94 | 85.6 | 82.06 | 86.4 | 85.18 | 85.92 | ||
| Flickr30k | 68.50 | 99.11 | 98.60 | 97.11 | 89.62 | 99.00 | 98.9 | ||
| Fused- Modal | TIT | WebQA | 54.84 | 78.42 | 32.42 | 66.99 | 49.62 | 82.24 | 84.11 |
| EDIS | 36.78 | 54.09 | 53.01 | 41.26 | 49.62 | 68.10 | 77.40 | ||
| ITT | OVEN | 22.32 | 45.98 | 23.69 | 0.38 | 14.4 | 59.67 | 64.13 | |
| INFOSEEK | 18.53 | 27.58 | 20.05 | 3.06 | 12.69 | 39.22 | 34.67 | ||
| ReMuQ | 76.20 | 83.71 | 26.41 | 94.60 | 52.15 | 96.73 | 95.48 | ||
| OKVQA | 17.14 | 17.44 | 9.67 | 13.28 | 16.71 | 30.08 | 32.61 | ||
| LLaVA | 72.81 | 91.91 | 51.64 | 53.18 | 77.48 | 98.93 | 98.18 | ||
| ITI | FashionIQ | 3.28 | 24.54 | 2.93 | 9.81 | 3.73 | 26.34 | 29.89 | |
| CIRR | 14.65 | 45.25 | 10.53 | 36.52 | 13.19 | 47.70 | 51.79 | ||
| ITIT | OVEN | 27.77 | 68.83 | 30.56 | 0.39 | 54.46 | 78.96 | 83.05 | |
| EVQA | 28.75 | 40.08 | 16.64 | 15.34 | 26.39 | 77.32 | 79.88 | ||
| INFOSEEK | 22.27 | 49.05 | 23.32 | 5.96 | 39.69 | 41.14 | 31.58 | ||
| Avg. | 37.32 | 43.66 | 42.01 | 50.04 | 42.52 | 63.42 | 65.87 |
8.1 Detailed Results on UMRB
@@
Table 7 presents the detailed evaluation results of the baseline systems alongside our GME on UMRB tasks. First, focusing on the average scores, our smaller model, i.e. GME-Qwen2-VL-2B, already outperforms the previous state-of-the-art UMR model (VISTA). The larger model, i.e. GME-Qwen2-VL-7B, further enhances this performance. In addition, focusing on specific scores on different datasets, our GME achieves state-of-the-art performance on each dataset except the Nights dataset. VISTA and CLIP-SF scored highly on the Nights dataset, likely due to their use of independent image and text encoders for cross-modal retrieval. In the II task, these models relied solely on the image encoder for encoding without cross-modal alignment, which may explain their superior performance on the Nights dataset.
| Type | Task | Dataset | TT | II | TVD | TI | ITIT | Mix |
| Single- Modal | TT | Arguan | 56.25 | 43.51 | 56.73 | 33.53 | 53.22 | 56.22 |
| NFCorpus | 35.23 | 28.89 | 33.23 | 33.18 | 30.48 | 35.76 | ||
| Quora | 87.82 | 74.37 | 86.32 | 86.43 | 85.2 | 87.4 | ||
| SCIDOCS | 19.07 | 11.82 | 17.51 | 17.2 | 16.93 | 19.88 | ||
| TRECCOVID | 75.57 | 47.89 | 50.89 | 72.37 | 58.92 | 76.38 | ||
| II | Nights | 27.97 | 28.11 | 24.9 | 28.53 | 26.04 | 30.85 | |
| Cross- Modal | TI | MSCOCO | 59.7 | 59.41 | 63.67 | 76.91 | 44.97 | 75.3 |
| Flickr30k | 83.92 | 65.52 | 87.32 | 93.18 | 74.52 | 93.06 | ||
| TVD | DocVQA | 35.8 | 24.24 | 48.38 | 40.58 | 28.05 | 45.62 | |
| Shift Project | 57.86 | 45.47 | 77.08 | 50.36 | 53.12 | 74.84 | ||
| IT | MSCOCO | 74.72 | 63.82 | 80.46 | 84.64 | 70.48 | 84.24 | |
| Flickr30k | 94.1 | 82.5 | 96.3 | 97.2 | 90.1 | 97.5 | ||
| Fused- Modal | ITT | LLaVA | 92.75 | 89.05 | 86.02 | 89.24 | 88.73 | 95.02 |
| ReMuQ | 89.61 | 85.47 | 76.45 | 85.12 | 86.73 | 89.75 | ||
| OKVQA | 24.55 | 16.6 | 15.78 | 16.92 | 18.57 | 20.23 | ||
| ITI | FashionIQ | 5.53 | 4.2 | 5.43 | 8.86 | 11.08 | 11.89 | |
| CIRR | 17.24 | 15.04 | 15.42 | 17.5 | 25.71 | 29.86 | ||
| ITIT | OVEN | 59.81 | 38.42 | 57.31 | 56.69 | 65.08 | 63.04 | |
@@
8.2 Detailed Results on UMRB-Partial
Figure 3 of main paper illustrates our exploration of the training data, as discussed in Section 4.2, with specific results presented in Table 8. This table details the scores of our models trained on six data types: TT, II, TVD, TI, ITIT, and Mix across various tasks. We find that the model trained on mixed data performs the best.
@@ 8.2 Detailed Results on UMRB-Partial
| GME-Qwen2-VL-2B | 53.31 | 61.52 | 42.3 | 38.13 | 46.31 | 92.6 | 45.3 | 72.93 | 40.88 | 37.2 | 60.01 | 87.24 | 23.17 | 63.82 | 29.06 | 59.24 | ||||||||||||||||||||||||||
| GME-Qwen2-VL-7B | 55.68 | 64.60 | 45.38 | 41.66 | 50.78 | 94.27 | 57.14 | 79.21 | 42.38 | 38.40 | 67.74 | 88.05 | 27.38 | 62.31 | 23.26 | 52.6 | ||||||||||||||||||||||||||
8.3 Detailed Results on BEIR
@@
| GME-Qwen2-VL-2B | 53.54 | 38.85 | 71.82 | 25.83 | 95.19 | 70.32 | 83.15 | 38.32 | 84.12 | 27.57 | 29.86 | 58.17 | 39.06 | 27.5 | 46.83 | 75.98 | 44.21 | |||||||||||||||||||||
| GME-Qwen2-VL-7B | 54.50 | 46.54 | 75.14 | 31.82 | 95.85 | 77.29 | 84.59 | 45.54 | 64.90 | 34.20 | 31.89 | 63.41 | 43.14 | 31.43 | 53.69 | 80.30 | 58.80 | |||||||||||||||||||||
8.4 Detailed Results on M-BEIR
M-BEIR, a multimodal benchmark for IR, serves as a comprehensive large-scale retrieval benchmark designed to evaluate multimodal retrieval models. As shown in Table 10, we report Recall@10 scores for the Fashion200K and FashionIQ datasets, while Recall@5 scores are provided for all other datasets. In M-BEIR, our GME continues to demonstrate state-of-the-art performance, underscoring the effectiveness of our approach.
@@ 8.4 Detailed Results on M-BEIR
| GME-Qwen2-VL-2B | 83.91 | 54.57 | 91.11 | 94.61 | 71.05 | 94.29 | 99.02 | 93.15 | 97.89 | 98.89 | 87.84 |
| GME-Qwen2-VL-7B | 87.58 | 56.63 | 92.39 | 94.58 | 76.12 | 97.26 | 99.63 | 95.89 | 99.5 | 99.63 | 89.92 |
8.5 Detailed Results on ViDoRe
ViDoRe represents the Visual Document Retrieval Benchmark, encompassing various page-level screenshot retrieval tasks. This benchmark includes the TVD tasks within our UMRB. Table 11 presents the detailed nDCG@5 scores for our GME and other models. Our smaller model, i.e. GME-Qwen2-VL-2B, surpasses the previous state-of-the-art model (ColPali), which was exclusively trained on this dataset for this specific task. The larger model, i.e. GME-Qwen2-VL-7B, further improves upon this performance.
9 Experiment Details
@@
9.1 Training Details
Our GME models (both 2B and 7B) are initialized using the Qwen2-VL [65] model series. We employ the transformers library for training in BF16 precision. The training utilizes Low-Rank Adaptation (LoRA) [19] with a rank of 8. We apply a decoupled AdamW optimizer with a learning rate and a weight decay of 1e-4. Additional hyperparameters are detailed in Table 12.